OpenStudy (anonymous):

The diameter of a circle with a circumference of 11.5 cm is ___? cm

Parth (parthkohli):

\( \color{Black}{\Rightarrow \large 2\pi{d \over 2} = 11.5 }\) Solve for \(d\).

OpenStudy (anonymous):

\(C=\pi d\)

OpenStudy (anonymous):

is 11.5 the solution? and how do i solve for d?

OpenStudy (anonymous):

nvm the first question

OpenStudy (anonymous):

hello?

jhonyy9 (jhonyy9):

C=2pir =11,5 diameter=2r so than 2pir=11,5 so 2rpi=11,5 so D*pi=11,5 so D=11,5/pi

OpenStudy (anonymous):

huh? so i just do 11.5(pi) ? @jhonyy9

jhonyy9 (jhonyy9):

yes good luck bye

OpenStudy (anonymous):

rude

OpenStudy (anonymous):

\[\Large C=\pi d \text{ implies } d=\frac{C}{\pi}\]